• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Huber

    Log entry time 23:44:31 on April 8,2001

    Entry number 60051

    keyword=Evening shift summary

    16:00 - Shift begins with first run after Hall C energy change.
    17:02 - Beam off because of RF trip. Beam back at 17:10
    17:26 - Start pulsed beam after Hall C entry. Then, beam trips off due to power supply communication problems. Waiting for an expert to arrive...

    22:30 - Right HV appears to have tripped off. At this time also noticed that we could not telnet into the port server, hatsv3. Jack Segal enters the hut to do a reset. See his message #60050.

    22:53 - Start pulsed beam after ARC power supply problems. In the process of checking out the orbits "the beam to Hall A looked really great", they lost the entire South Linac RF. I guess this will take a while to reestablish too...

    23:45 - Incoming MCC crew chief calls to say that it will be at least 2 hours before the RF problems are rectified and beam back.


    Run 3394 Start=15:50 Stop=17:02 Events=830k
    Run 3395 Start=17:10 Stop=17:44 Events=182k